myininaya (myininaya):

So remember that \[ \lim_{ n \rightarrow \infty} \frac{1}{n^2}=0 .\] Why is it that you can't say the limit for the following is 0? \[\lim_{n \rightarrow \infty } \frac{1+2+ \cdots +n}{n^2}=0+0+0+ \cdots +0=0\] but this limit is suppose to be 1/2 like this is probably a stupid question

OpenStudy (anonymous):

How come now you have:\[ \lim_{n\to 0}\frac 1{n^2} =0 \]

myininaya (myininaya):

oops that was suppose to be infinity there at that one spot

myininaya (myininaya):

no this just a question i have been thinking about for awhile

OpenStudy (anonymous):

The reason you can't do it is because the rule for limits over addition can work a finite number of times, but not an infinite number of times.

OpenStudy (anonymous):

Recall that \(1+2+\cdots+n = \dfrac{n(n+1)}{2}\). So, \(\displaystyle \lim_{n\to 0}\frac{1+2+\cdots n}{n^2 } = \lim_{n\to0} \frac{n(n+1)}{2n^2} = \frac{1}{2}\lim_{n\to 0} \left(1+\frac{1}{n}\right)\). However, this doesn't have a limit as \(x\to0\). It has a limit of \(\dfrac{1}{2}\) as \(x\to\infty\), though. Is that what you meant? :-)

OpenStudy (anonymous):

I mean as \(n\to 0\), not \(x\to 0\).

OpenStudy (anonymous):

...and the same for \(n\to\infty\) instead of \(x\to\infty\) (I'm getting my variables mixed up for some reason)

OpenStudy (anonymous):

"The Limit Does not Exist" - Mean Girls

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}(\frac{1+2+\cdots +n}{n^2})=\lim_{n \rightarrow \infty}(\frac{1}{n^2})+\lim_{n \rightarrow \infty}(\frac{2}{n^2})+\cdots +\lim_{n \rightarrow \infty}(\frac{n}{n^2}) \\=0+0+\cdots+0=0\] so by what wio said we can't do this

myininaya (myininaya):

And I know that 1+2+...+n=n(n+1)/2

OpenStudy (anonymous):

Basically, I would suggest that: \[ \lim_{n\to \infty}\frac{\sum_{k=1}^n k}{n^2} \neq \sum_{k=1}^n\lim_{n\to \infty}\frac k{n^2} \]

OpenStudy (anonymous):

Well, I am being a bit presumptuous. Basically the problem lies somewhere before the first step and the second step.

myininaya (myininaya):

Ok thanks you guys I guess I just forgot that we can only do that addition thing a finite number of times That makes sense wio

ganeshie8 (ganeshie8):

.....99999999999999999999999999999 +1 ---------------------------------------- 0

ganeshie8 (ganeshie8):

for some reason ur infinite additions reminds of that hmm

myininaya (myininaya):

so smart

OpenStudy (anonymous):

It's a good example of a case where induction doesn't work to infinity.

myininaya (myininaya):

I blame me forgetting this because of that feral cat that I mistook as a rat. I went outside and I heard something in a box. Ran back inside. But it was just a kitty. Small and cute.

myininaya (myininaya):

I was kidding. I just forgot. That is all.

myininaya (myininaya):

But there really was a kitty outside today that I thought was a rat.

OpenStudy (ikram002p):

at some point it will be recursive like 1/n^2+2/n^2+3/n^2+....+(n-3)/n^2+(n-2)/n^2+(n-1)/n^2+n/n^2 1/n^2+2/n^2+3/n^2+....-3/n-2/n-2/n -1/n +n/n so u would have sum i /n^2 as n goes to infinity - sum i /n as n goes to infinity + 1 :O that leads to nothing :3 but just away to say not all terms goes to 0